For a recent semester, the number of men enrolled in elementary algebra totaled 7 more than 4 times the number of women. Find the number of men enrolled in elementary algebra given that the total number of students enrolled was 187.
Answer by ikleyn(52908)

.

men   + women = 187


(4x+7) + x    = 187


5x = 187 - 7 = 180


 x           = 180/5 = 36,  number of women.


ANSWER.  187 - 36 = 151 men.


CHECK.  151 = 4*36 + 7 = 144 + 7.    ! Correct !

Solved.